OpenStudy (anonymous):

FAN AND MEDAL I just need help on this worksheet. It's super easy but I suck at math. I have answered 1-5 (I got A on all of them, which I also am not 100% sure if that is correct.) and I need help on the rest. I will fan and medal just for helping. file:///C:/Users/Hannah's%20Laptop/Downloads/Algebra%202%202nd%20Semester%20Unit%201%20Portfolio%20(2).pdf

OpenStudy (anonymous):

where r the questions?

OpenStudy (anonymous):

File://C: says it is still on your hard-drive!

OpenStudy (anonymous):

Oh gosh, I didn't even see that! This should be it

OpenStudy (anonymous):

you again? wow

OpenStudy (anonymous):

me again!

OpenStudy (anonymous):

like black american history?

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

you must forgive me, most mathematicians have memories like this!

OpenStudy (anonymous):

well, at least you know I love Rosa Parks, right?

OpenStudy (anonymous):

Oh you are asking me if I like Black History! haha, yes, I find it very interesting. I am one of the few people who really enjoys learning. But I hate school and try to whiz past it and find any shortcut possible at the same time, I enjoy learning by my own methods.

OpenStudy (anonymous):

So do you know how to do the problems in the worksheet after number 5?

OpenStudy (anonymous):

hey, I am a tutor who charges 30 euro per hour for my time, so do you have a question ?

OpenStudy (anonymous):

Wow, that's pretty dang expensive! Haha. Yes, the file I attached previously is a worksheet I need help with on questions after 5.

OpenStudy (anonymous):

wait a minute

OpenStudy (anonymous):

the problem is you gave your hard-drive adress

OpenStudy (anonymous):

are we supposed to break into your hard-drive?

OpenStudy (anonymous):

I reposted a file attached to a comment earlier, I am not sure if you saw it but I will attach it again

OpenStudy (anonymous):

Does that work?

OpenStudy (anonymous):

break the first question into factors, that ius how to solve it

OpenStudy (anonymous):

I am referring to questions 5-13. By first question do you mean 5?

OpenStudy (anonymous):

so you have solved questions 1-5?

OpenStudy (anonymous):

Yes, that is correct

OpenStudy (anonymous):

Am I right you ned question 6 solved?

OpenStudy (anonymous):

(2-x-1)/(3-x)

OpenStudy (anonymous):

you do know this a terrible question!

OpenStudy (anonymous):

Yes I do! haha

OpenStudy (anonymous):

well, the good news is the -x's cancel so you are left with x= 1/3

OpenStudy (anonymous):

I am beginning to think the answer to all of these are A. ;)

OpenStudy (anonymous):

what about 7?

OpenStudy (anonymous):

first thing is cross multiply! 3(x-1)=2(x+1)

OpenStudy (anonymous):

3x-3 = 2x+2

OpenStudy (anonymous):

So 3x+3 and 2x_2

OpenStudy (anonymous):

right then haha

OpenStudy (anonymous):

you know babes!

OpenStudy (anonymous):

so then what?

OpenStudy (anonymous):

3x+3=2x+2 bring real numbers together and x's together this is what we call gathering terms babe! 3x(-2x)+3 = 2x (-2x) +2 x+3=2

OpenStudy (anonymous):

A again! Man. Okay, 8?

OpenStudy (anonymous):

right so, x+3 = 2 therefore x = 2-3 x= -1

OpenStudy (anonymous):

now you need to check it

OpenStudy (anonymous):

substitute x into original equation see if it equals zero

OpenStudy (anonymous):

One side equals negative two and one .3 repeating. I don't think so

OpenStudy (anonymous):

lets see what happened

OpenStudy (anonymous):

okay haha

OpenStudy (anonymous):

this is Q.7 yes?

OpenStudy (anonymous):

straight away -x cancels

OpenStudy (anonymous):

that leaves 2 -1 over 3

OpenStudy (anonymous):

Wait number seven would be 5, or D, because once you go through and plug in each answer for x, D is the only one that works.

OpenStudy (anonymous):

that meaans your answer is -1/3

OpenStudy (anonymous):

test 2 - (-1/3)-1/3-(-1/3)

OpenStudy (anonymous):

2+ 1/3 -1/3(-1/3)

OpenStudy (anonymous):

Yup I'm lost.

OpenStudy (anonymous):

dont worry will explain

OpenStudy (anonymous):

okay, haha

OpenStudy (anonymous):

sorry it's 1AM here and my eyes are hardly open. I am basically useless right now.

OpenStudy (anonymous):

Actually I should probably go to bed now, I have to wake up very early to take my brother to school, can we talk more tomorrow?

OpenStudy (anonymous):

yes, the answer is 21/30

OpenStudy (anonymous):

Thanks! I will come back around again tomorrow. /night